OpenStudy (anonymous):

24 out of 240 A.1% B.10%

OpenStudy (anonymous):

i will give you a medal if you answer

OpenStudy (anonymous):

B

OpenStudy (kendricklamar2014):

Dont give direct answers @borak

OpenStudy (anonymous):

to know the rate you divide the 24 on 240; 24/240 = 1/10 and = 10%

OpenStudy (anonymous):

ok ty

OpenStudy (kendricklamar2014):

DO u have a calculator @Alison124

OpenStudy (anonymous):

yes i do